OpenStudy (anonymous):

What is the range of the graph of y = 3(x – 3)2 + 4?

OpenStudy (anonymous):

what does range mean?

OpenStudy (anonymous):

slope

OpenStudy (icalibear):

what is the slope?

OpenStudy (anonymous):

herere the answers y less than or greater to 4 y greater than or equal to 4 y greater than or equal to 3 y less than or greater to 3

OpenStudy (anonymous):

3x - 9(2) +4 3x - 18 + 4 3x + 22 (0, 22) and (1, 7) 22 - 7 = 15 15/-1 -15 is the slope

OpenStudy (icalibear):

so it's the last one isnt it?

OpenStudy (anonymous):

icalibear: algebra i

OpenStudy (icalibear):

yeah ill let you help with this one im doing geometry at the moment my mind is somewhere else lol

OpenStudy (anonymous):

\[y = 3(x – 3)^2 + 4\]is this the equation?

OpenStudy (anonymous):

oh if it's that my answer is wrong I'll redo it

OpenStudy (anonymous):

\[y = 3(x – 3)^2 + 4\]\[y=3(x^2-6x+9)+4\]\[y=3x^2-18x+27+4\]\[y=3x^2-18x+31\]\[y'=6x-18\]That's the slope. I don't know what range means

OpenStudy (anonymous):

Range is all the y values of a graph so the range is \[y \ge 4\] because this a parabola opening upwards with vertex at (3,4)